Father is five times as old as his son now. In six years,he will be three times as old as his son. How old is he now?
Answer by macston(5194)

F=father's age; S=son's age
.
F=5S
.
F+6=3(S+6)
5S+6=3S+18
2S=12
S=6 ANSWER 1: Son is 6 years old.
.
F=5S
F=5(6)=30 ANSWER 2: Father is 30 years old.
.
CHECK:
F+6=3(S+6)
30+6=3(6+6)
36=3(12)
36=36